# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v2

@@ 5.2.1 Model Selection: Conventional Machine Learning or Deep Learning?
In PHM data challenge competitions, participants face the critical decision between employing conventional ML algorithms or opting for advanced DL models. This choice is influenced by various factors: the volume and quality of available data, and the complexity or nature of the research task (whether it be classification or prediction). An analysis of research papers from the past six years reveals insightful trends and preferences in model selection:
@@
This trend is clear when addressing classification problems within PHM competitions. As depicted in Table 2, Table 5, and Figure 2, 2022 PHM NA and 2023 IEEE competitions, because of their large datasets, have facilitated the application of DL techniques. Specifically, the 2023 IEEE competition saw widespread use of deep CNN models, including ensemble-based CNNs [62] and residual-based CNNs [91, 51], showcasing advanced DL’s capability to tackle complex problems. Additionally, 2022 PHM NA competition favored diverse DL approaches like metric learning + pseudo label-based DL [80], CNN + DANN [49], X-Vectors [65], indicating a preference for integrating transfer learning with DL for enhanced performance. However, the scenario differed for the 2022 PHM EU competition, which provided approximately 2000 samples and encountered data imbalance challenges. Given these constraints, many teams leaned towards tree-based methods known for their robustness and ability to address data imbalance, deploying algorithms like XGBoost [32, 102], LightGBM [102], decision trees [75], and random forest [104, 75]. In competitions with significantly fewer data samples, such as the 2021 PHM EU and 2023 PHM AP, where the dataset sizes were around 100-200 samples, the deployment of DL was impractical due to its requirement for a large volume of data. Instead, conventional ML methods proved more effective. For the 2021 PHM EU, tree-based methods were predominantly used [22, 41, 4, 85, 106], while the 2023 PHM AP competition saw a variety of approaches, including similarity-based [73, 47], rule-based [61], KNN/K-means [47, 2], and tree-based methods [61, 2], emphasizing the adaptability of traditional methods to limited data scenarios.
@@
In the realm of RUL prediction tasks, the inherent long-time series nature of the datasets, even when datasets are smaller (less than 100 samples), allows for different approaches to data utilization and data augmentation. By considering each time point or a sequence of time points (data window) as an independent training sample, the effective size of the dataset can be substantially increased. This is evident from Table 7 and Figure 2, which indicate the application of both DL and conventional ML methods across various competitions, with the notable exception of the 2019 PHM NA competition, where only conventional ML methods were employed. As detailed in Figure 2, competitions such as 2018 PHM NA, 2020 PHM EU, and 2021 PHM NA saw widespread adoption of diverse DL techniques, ranging from CNN [113, 97, 24], RNN-based [38, 118, 107] models to transfer learning [66, 107], Autoencoder-based [42], and transformer-based [124, 68] methods, to tackle RUL prediction challenges. Conversely, the 2019 PHM NA competition’s unique challenge, due to the partial absence of wave data in the test dataset, necessitated a hybrid approach that integrated data-driven (conventional ML) and physics-based approaches [50, 121] to address the problem.
(2) The complexity or nature of the research task influences the choice of modeling approach to some degree. Figure 3 shows that tree-based methods (such as Decision Trees, Random Forest, and XGBoost) are favored in conventional ML and DL solutions often favor LSTM-based and CNN-based approaches. A closer examination of Figure 2, Table 4, and Table 7 reveals key insights. The choice between conventional ML and DL for detection and diagnosis (classification problem) hinges largely on dataset size, as discussed before. Absent this factor, for the RUL prediction problem, the majority of teams opted for DL approaches. This is because RNN-based methods (LSTM, GRU, etc) are proficient in handling sequential, time-series data, and CNN-based methods are adept at feature extraction from multidimensional datasets, enhancing the capability to tackle RUL prediction problems efficiently. While conventional ML methods have also shown efficacy, the scalability of DL indicates its superior potential for RUL prediction, given sufficient data. When domain shift is present in the competition dataset, transfer learning and domain adaptation methods are often proposed to adjust the model to new data distributions, and address limitations in training data. This adaptability ensures the model’s applicability and effectiveness under varied operational conditions, a frequent occurrence in PHM scenarios.
@@
(3) When datasets become publicly accessible, there is a clear shift towards the adoption of more sophisticated DL approaches in PHM data challenge competitions over time. Taking the 2018 PHM NA competition as a case study, our review of 12 published papers utilizing this dataset reveals that the methodologies initially favored by the competition teams largely comprised tree-based methods, SVMs, and basic LSTM algorithms [38, 93]. However, as time goes by, there is a discernible trend towards the development of more complex DL algorithms. This includes but is not limited to, GRU [118], knowledge distillation [126], a two-stage deep transfer learning framework utilizing TCN and DANN [66], TCN combined with LSTM [35], attention mechanisms [35, 68], and transformers [124, 68]. This trend underscores a growing need for advancing DL approaches in the PHM domain, to deepen dataset understanding and enhance model performance. The continuous innovation and refinement of DL methods indicate a vibrant future for PHM, where advanced data-driven analyses pave the way for novel discoveries in the field.
(4) A general discussion on the advantages and limitations of CNN, RNN, and tree-based methods in the context of PHM reveals: (1) CNNs excel in extracting features from multidimensional data but struggle with interpretability and require substantial computational resources. (2) RNNs are adept at handling sequential time-series data, benefiting from their ability to remember information over long sequences, useful for understanding long-term dependencies. These approaches are helpful for health state monitoring and RUL prediction, though their complexity and data demands can be challenging. (3) Tree-based methods are valued for their interpretability, feature importance analysis, ability to handle various data types, and effectiveness in capturing non-linear relationships, but they might not achieve the neural networks’ accuracy on large complex datasets.
In summary, selecting an appropriate ML approach should consider the dataset’s specific characteristics, aiming to balance accuracy, interpretability, and computational efficiency to tackle PHM’s diverse challenges. When training data are sufficient, DL methods typically offer superior performance compared to conventional ML approaches.
@@ 5.2.2 Machine Learning Model Interpretability
In PHM data challenge competitions, accuracy in prediction and classification is still the primary goal. However, with the advance of ML and DL, emphasizing the importance of model interpretability is becoming as crucial as their accuracy in prediction and classification due to its significant impact on safety, operational decision-making, and compliance with industry standards. Model interpretability not only fosters trust and facilitates collaboration between data scientists and domain experts, but it also plays a pivotal role in error analysis and model refinement in PHM.
While only a few teams in competitions have used model interpretability methods to explain the output of their models, outside of competitions, various techniques have been developed and applied to enhance ML model interpretability [64, 6]. As the complexity of deep neural network models, often referred to as ”black boxes”, some previous research utilized interpretable methods like Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping in CNNs, and attention mechanisms in sequential models to shed light on model decision-making processes [98]. Additionally, model-agnostic methods like LIME [88] and SHAP [70, 101] have been instrumental in offering insights into model behavior. For instance, Baptista et al. applied the SHAP model to evaluate the outcomes of three different algorithms (Linear Regression, MLP, and Echo State Network) using the Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset (jet engines) [9]. Moreover, Moradi et al. introduced an interpretable artificial neural network designed for the automatic selection and fusion of features to develop optimal health indicators from data gathered through structural health monitoring (SHM) [77]. From a broader prospect, the challenge in PHM is to balance the trade-off between model accuracy and interpretability. Achieving this balance is not just about enhancing model performance, but also about building trust and ensuring compliance with industry standards. Furthermore, in an era increasingly focused on ethical and responsible AI, transparent and interpretable models are key to not only enhancing the technical aspects of AI solutions in PHM but also extending to ensuring their successful integration and acceptance in real-world applications [112].
@@ 5.2.3 Model Robustness and Generalization
In the realm of PHM, model robustness and model generalization are critical factors that determine the effectiveness of a model in real-world applications. Model robustness is defined as the ability of a model to maintain its performance when facing diverse challenges like noise, outliers, and adversarial examples. Techniques such as data augmentation, where training data can be expanded by creating modified versions of existing data or synthesizing new data, have become commonplace. For instance, in the 2023 IEEE competition, Kreuzer’s team employed methods like additive white Gaussian noise, circular shift, and random amplitude scaling to increase the volume of training data [51]. Moreover, regularization techniques like L1 and L2 regularization have been instrumental in preventing overfitting and enhancing stability, making the model less sensitive to small fluctuations in input data. What’s more, Ensemble methods have been increasingly recognized for their significant contribution to model robustness in PHM. A comprehensive analysis of recent studies shows that out of sixty research papers, nine utilized various ensemble techniques - including decision trees, random forest, XGBoost, ensemble LSTM [7], soft voting ensemble [49], transfer ensemble learning [107], ensemble regression [86], and ensemble CNN-based [62] approaches - across six different PHM data challenge competitions. This trend underscores the effectiveness of ensemble methods in bolstering model robustness, particularly by reducing the influence of noise and outliers, thereby enhancing the reliability of predictive models in PHM applications. Additionally, adversarial training, which involves training models on both regular and adversarial data, has been recognized for its potential to fortify models against adversarial attacks [117, 49, 83].
@@
Alongside robustness, model generalization has also been a focal point in PHM, with the goal being to develop models that perform reliably on new, unseen data. One key technique for improving generalization is cross-validation, which involves dividing the training data into several subsets and each time using one of the subsets as the validation data and others as the training data for model validation during the training process. In a notable example from the 2021 PHM NA competition, researchers opted for k-fold repeated random subsampling validation. This approach was chosen as an alternative to standard k-fold cross-validation to address its limitation, wherein the size of the validation set diminishes as the number of folds (k) increases [97]. In addition to cross-validation, transfer learning and domain adaptation methods have been crucial in maintaining model effectiveness and generalization across varied data distributions. These methods are particularly relevant in scenarios characterized by ”domain shift,” where the data from the target domain differs from the source domain. The specific implementations of these methodologies in PHM competitions have been detailed in the earlier subsubsection titled “Domain Shift”. Outside of PHM data challenges, some novel approaches are proposed to deal with model generalization issues. Matthew Russell and Peng Wang adopted a domain adversarial transfer learning method inspired by generative adversarial networks, utilizing a 1D CNN architecture to predict tool wear on unseen domains using NASA’s milling dataset [115]. Ding et al. developed a multi-source domain generalization learning approach (GRU + Transformer) that can effectively learn useful degradation feature representations from various run-to-failure datasets of internal combustion engine journal bearings across different conditions and predict unseen working conditions well [26]. Furthermore, Ding et al. proposed an adversarial out-domain augmentation (AOA) framework for predicting the RUL of bearings under unseen conditions. The effectiveness of this AOA-based RUL prediction was validated using IEEE PHM Challenge 2012 and XJTU-SY run-to-failure datasets, illustrating its robustness in domain generalization for predictive maintenance [27].
@@ 5.3.1 A Lack of Multi-Modal Machine Learning Analysis
Our review of PHM data challenge competitions reveals a reliance on single-modality data in the competitions, such as pressure signals, currents, vibrations, or images, without incorporating multi-modal datasets for fault diagnosis and prognosis. Multi-modal machine learning (MMML) in PHM refers to capturing complementary information from multiple sources (different data types) to achieve a more comprehensive and precise evaluation of PHM tasks [84, 109]. Despite its potential, it is still underexplored in the PHM domain.
Recently, Jiang et al. leveraged two modality data (vibration and current signals) to develop deep belief networks (DBNs) for diagnosing wind turbine gearbox faults [45]. Su et al. proposed an MMML model using parametric specifications, text descriptions, and images of vehicles to predict five vehicle rating scores [100]. Additionally, Fan et al. evaluate several MMML strategies to create a comprehensive PHM system for coolant pumps in commercial heavy-duty vehicles, utilizing data from onboard signals, multi-dimensional histograms, and categorical variables [31]. Wang et al. proposed a novel method for feature fusion in multimodal data (vibration and torque signals), applying it to diagnose the bearings faults [114].
Literature on MMML for PHM tasks shows that most of the methods are focused on classification problems. Therefore, there is still room for further exploration of the application of MMML to prognosis challenges. Moreover, at the heart of MMML for PHM tasks is getting useful multi-modal latent representations using effective data fusion techniques. While simple concatenation methods are widely used for data fusion in MMML, there’s a need for further exploration of more advanced fusion techniques, such as attention-based or transformer-based mechanisms. For future PHM data challenges, the provision of open-source multi-modal datasets would empower researchers to investigate and apply more advanced MMML techniques, potentially leading to significant advancements in the context of PHM.
@@ 5.3.2 A Lack of Adopting Systematic Approaches for Effective PHM Systems Construction
Our analysis of ML and DL methods across nine open-source industrial datasets has revealed the transformative potential that ML approaches can have on PHM. Nevertheless, the nature of the competition tends to prioritize solutions that chiefly enhance accuracy, potentially at the expense of systematic approach, reusability, and methodological inheritance. It is therefore vital to pursue systematic methodologies for constructing effective PHM systems that go beyond the competitive framework. This should involve thorough research and analysis of open-source datasets to advance ML and DL strategies, aiming not just for competition success but also for benchmarking and comparative analysis.
Souza et al. devised an ML-based, data-oriented pipeline for constructing a Prognosis and Health Management System (PHMS) focused on RUL prediction, utilizing semi-supervised ML with Autoencoder, XGBoost, and the SHAP method [99]. Professor Enrico Zio highlighted the primary challenges and future directions for implementing condition-based and predictive maintenance in real-world applications [127]. As shown in Figure 5, Hu et al. offered a new perspective on reviewing PHM efforts by proposing a division of the PHM lifecycle into DEsign, DEvelopment, and DEcision ($DE^{3}$) phases, and showcasing the pivotal activities and challenges within these stages [36]. Additionally, Lee et al. introduced a novel SoQ methodology for multi-stage manufacturing processes. It can help to analyze multi-parameter influences on product quality and model inter-process relationships in multi-stage manufacturing systems [58]. Additionally, Figure 6 shows an Industrial Large Knowledge Model framework proposed by Lee’s team. This framework aimed to utilize the power of LLMs and domain-specific knowledge to address the complex challenges in the PHM domain [59].
Moving forward, developing novel, systematic approaches for PHM systems should be encouraged in future PHM data challenge competitions. These efforts can notably augment the systematization and applicability of ML and DL approaches, thereby expanding their utility beyond the confines of the competition-centric paradigm. Such advancements promise to narrow the divide between academic research and industrial practice, facilitating the broader adoption of data-driven ML across various industrial contexts.

### v3

@@ 4.2.1 Model Selection: Conventional Machine Learning or Deep Learning?
In PHM data challenge competitions, participants often grapple with the difficult decision of whether to use conventional ML algorithms or advanced DL models. This choice is influenced by various factors: the volume and quality of available data, and a trade-off between computational cost and performance. An analysis of research papers from the past six years reveals insightful trends and preferences in model selection:
@@
This trend is clear when addressing classification problems within PHM competitions. As depicted in Table 4 and Figure 2, 2022 PHM NA and 2023 IEEE competitions, because of their large datasets, have facilitated the application of DL techniques. In 2023 IEEE, various deep CNN models are proposed, including ensemble-based CNNs [66] and residual-based CNNs [97, 55]. Additionally, in 2022 PHM NA, many teams integrate diverse DL approaches with transfer learning such as metric learning + pseudo label-based DL [86], CNN + DANN [53], X-Vectors [69]. However, the scenario differed for the 2022 PHM EU competition, which provided approximately 2000 samples and encountered data imbalance challenges. Given these constraints, many teams leaned towards tree-based methods known for their robustness and ability to address data imbalance, deploying algorithms like XGBoost [32, 108], LightGBM [108], decision trees [80], and random forest [110, 80]. In competitions with significantly fewer data samples, such as 2021 PHM EU and 2023 PHM AP, where the dataset sizes were around 100-200 samples, the deployment of DL was impractical due to its requirement for a large volume of data. Instead, conventional ML methods proved more effective. In 2021 PHM EU, tree-based methods were predominantly used [23, 45, 4], while in 2023 PHM AP, similarity-based [78, 51], rule-based [65], KNN/K-means [51, 2], and tree-based methods [65, 2] are proposed, emphasizing the adaptability of traditional methods to limited data scenarios.
@@
In the realm of RUL prediction tasks, the inherent long-time series nature of the datasets, even when datasets are smaller (less than 100 samples), allows for different approaches to data utilization and data augmentation. By considering each time point or a sequence of time points (data window) as an independent training sample, the effective size of the dataset can be substantially increased. This is evident from Table 6 and Figure 2, which indicate the application of both DL and conventional ML methods across various competitions such as CNN-based [120, 103, 25], RNN-based [125, 113], transfer learning [70, 113], Autoencoder-based [46], transformer-based [129, 72] methods, and conventional ML with physics-based approaches [54, 127] to tackle RUL prediction challenges.
(2) When selecting DL models, there is often a trade-off between computational cost and performance. In the competitions, participants may choose highly complex DL models that require substantial computational resources to achieve even a small improvement in accuracy. While this approach might secure a higher position on the leaderboard, it may not be the most practical choice for real-world implementation. In practice, the value of such a small accuracy improvement needs to be weighed against the increased computational cost and potential scalability issues.
@@
(3) When datasets become publicly accessible, there is a clear shift towards the adoption of more sophisticated DL approaches in PHM data challenge competitions over time. Taking the 2018 PHM NA competition as a case study, our review of 12 published papers utilizing this dataset reveals that the methodologies initially favored by the competition teams largely comprised tree-based methods, SVMs, and basic LSTM algorithms [42, 99]. However, as time goes by, there is a discernible trend towards the development of more complex DL algorithms. This includes but is not limited to, GRU [125], knowledge distillation [131], a two-stage deep transfer learning framework utilizing TCN and DANN [70], TCN combined with LSTM [39], attention mechanisms [39, 72], and transformers [129, 72]. This trend underscores a growing need for continuous innovation and refinement of DL methods in the PHM domain.
@@ 4.2.2 Machine Learning Model Interpretability
In PHM data challenge competitions, accuracy in prediction and classification is still the primary goal. However, with the advance of ML and DL, emphasizing the importance of model interpretability is becoming as crucial as their accuracy in prediction and classification, because it helps to illustrate the model decision-making process and plays an important role in error analysis and further model refinement in PHM.
While only a few teams in competitions have used model interpretability methods to explain the output of their models, outside of competitions, various techniques have been developed and applied to enhance ML model interpretability [68, 6]. As the complexity of deep neural network models, often referred to as ”black boxes”, some previous research utilized interpretable methods like Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping in CNNs, and attention mechanisms in sequential models to shed light on model decision-making processes [104]. Additionally, model agnostic methods like SHAP [74, 107] and LIME [94]have been instrumental in offering insights into model behavior. For instance, Baptista et al. applied the SHAP model to evaluate the outcomes of three different algorithms (Linear Regression, MLP, and Echo State Network) using the Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset (jet engines) [9]. Moreover, Moradi et al. introduced an interpretable artificial neural network designed for the automatic selection and fusion of features to develop optimal health indicators from data gathered through structural health monitoring (SHM) [82]. Furthermore, in an era increasingly focused on ethical and responsible AI, transparent and interpretable models are key to not only enhancing the technical aspects of AI solutions in PHM but also extending to ensuring their successful integration and acceptance in real-world applications [119].
@@ 4.2.3 Model Robustness and Generalization
Model robustness is defined as the ability of a model to maintain its performance when facing diverse challenges like noise, outliers, and adversarial examples. Techniques such as data augmentation, where training data can be expanded by creating modified versions of existing data or synthesizing new data, have become commonplace. For instance, in 2023 IEEE competition, Kreuzer’s team employed methods like additive white Gaussian noise, circular shift, and random amplitude scaling to increase the volume of training data [55]. Moreover, regularization techniques like L1 and L2 regularization have been instrumental in preventing overfitting and enhancing stability, making the model less sensitive to small fluctuations in input data. What’s more, ensemble methods have been increasingly recognized for the contribution to model robustness in PHM. A comprehensive analysis of recent studies shows that out of 59 research papers, nine utilized various ensemble techniques - including tree-based methods, ensemble LSTM [7], soft voting ensemble [53], transfer ensemble learning [113], ensemble regression [92], and ensemble CNN-based [66] approaches - across six different PHM data challenge competitions. Additionally, adversarial training, which involves training models on both regular and adversarial data, has been recognized for its potential to fortify models against adversarial attacks [124, 53, 89]. Moreover, it is also important to consider probabilistic machine learning techniques, such as Bayesian networks, Gaussian processes, and probabilistic graphical models, which incorporate probability theory into the modeling process to handle uncertainty issues. They are crucial for dealing with uncertainties and improving the robustness of PHM models [33, 83, 36].
@@
Alongside robustness, model generalization refers to developing models that perform reliably on new, unseen data. One of the techniques for improving generalization is cross-validation, which involves dividing the training data into several subsets and each time using one of the subsets as the validation data and others as the training data for model validation during the training process. In 2021 PHM NA, researchers opted for k-fold repeated random subsampling validation to address its limitation, wherein the size of the validation set diminishes as the number of folds (k) increases [103]. In addition to cross-validation, transfer learning, and domain adaptation methods have been crucial in maintaining model effectiveness and generalization when data from the target domain differs from the source domain. The specific implementations in PHM competitions have been detailed in the earlier subsubsection titled “Domain Shift”. Outside of PHM data challenges, some novel approaches are proposed to deal with model generalization issues. Matthew Russell and Peng Wang adopted a domain adversarial transfer learning method inspired by generative adversarial networks, utilizing a 1D CNN architecture to predict tool wear on unseen domains using NASA’s milling dataset [122]. Ding et al. developed a multi-source domain generalization learning approach (GRU + Transformer) that can effectively learn useful degradation feature representations from various run-to-failure datasets of internal combustion engine journal bearings across different conditions and predict unseen working conditions well [27]. Furthermore, Ding et al. proposed an adversarial out-domain augmentation (AOA) framework for predicting the RUL of bearings under unseen conditions. The effectiveness of this AOA-based RUL prediction was validated using IEEE PHM Challenge 2012 and XJTU-SY run-to-failure datasets, illustrating its robustness in domain generalization for predictive maintenance [28].
@@ 4.3.1 A Lack of Multi-Modal Machine Learning Analysis
Our review of PHM data challenge competitions reveals a reliance on single-modality data in the competitions, such as pressure signals, currents, vibrations, or images, without incorporating multi-modal datasets for fault diagnosis and prognosis. Multi-modal machine learning (MMML) in PHM refers to capturing complementary information from multiple data sources (different types) to achieve a more comprehensive and precise evaluation of PHM tasks [90, 115]. Despite its potential, it is still underexplored in the PHM domain.
Recently, Jiang et al. leveraged two modality data (vibration and current signals) to develop deep belief networks (DBNs) for diagnosing wind turbine gearbox faults [49]. Su et al. proposed an MMML model using parametric specifications, text descriptions, and images of vehicles to predict five vehicle rating scores [106]. Additionally, Fan et al. evaluate several MMML strategies to create a comprehensive PHM system for coolant pumps in commercial heavy-duty vehicles, utilizing data from onboard signals, multi-dimensional histograms, and categorical variables [31]. Wang et al. proposed a novel method for feature fusion in multimodal data (vibration and torque signals), applying it to diagnose the bearings faults [121]. For future PHM data challenges, the provision of open-source multi-modal datasets would empower researchers to investigate and apply more advanced MMML techniques, potentially leading to advancements in the context of PHM.
@@ 4.3.2 A Lack of Adopting Systematic Approaches for Effective PHM Systems Construction
Our analysis of ML and DL methods across nine open-source industrial datasets has revealed the advantages of ML methods applied in PHM. Nevertheless, the nature of the competition tends to prioritize solutions that chiefly enhance accuracy, potentially at the expense of a systematic approach, reusability, and methodological inheritance. It is therefore vital to pursue systematic methodologies for constructing effective PHM systems that go beyond the competitive framework. This should involve thorough research and analysis of open-source datasets to advance ML and DL strategies, aiming not just for competition success but also for benchmarking and comparative analysis.
Souza et al. devised an ML-based, data-oriented pipeline for constructing a Prognosis and Health Management System (PHMS) focused on RUL prediction, utilizing semi-supervised ML with Autoencoder, XGBoost, and SHAP method [105]. As shown in Figure 5, Hu et al. offered a new perspective on reviewing PHM efforts by proposing a division of the PHM lifecycle into DEsign, DEvelopment, and DEcision ($DE^{3}$) phases, and showcasing the important activities and challenges within these stages [40]. Additionally, Lee et al. introduced a novel SoQ methodology for multi-stage manufacturing processes. It can help to analyze multi-parameter influences on product quality and model inter-process relationships in multi-stage manufacturing systems [62]. Moving forward, developing novel, systematic approaches for PHM systems should be encouraged in future PHM data challenge competitions. These efforts can augment the systematization and applicability of ML and DL approaches, thereby expanding their utility beyond the confines of the competition-centric paradigm. Such advancements promise to narrow the divide between academic research and industrial practice, facilitating the broader adoption of data-driven ML across various industrial contexts.
Figure 6: Industrial Large Knowledge Model Framework [63]